Shift report 20030118 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 18 2003 - 16:25:40 EST

  • Next message: Apache: "Shift report 20030118 16:00-24:00"
    ------------------------------------------------------------
    Date: Sat Jan 18 16:25:39 2003, Shift: 08:00-16:00
    Supervisor: cr, with: 
    Log book#13 pages 86 - 89
    
    Shift Summary:
    ------------------------------------------------------------
    1. RHIC Controlled Access in the morning. Since 1 pm we had good beam but some problems with detectors. Right now the beam is for Phenix background studies, with a huge background.
    
    2. Runs 7461 - 7474 all debugging runs, with different scaledown for triggers and RunError, which determined that the runs could not reach the DB. 
    
    3. During this shift, RD, JIJ and RP worked in IR. See logbook and E-log for summary. T3-map was often signaled by DAQ. JJG is debugging right now the DAQ. TPM1, TPM2 had high current and were turn to parking mode.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Jan 18 2003 - 16:29:30 EST